A reversible power cycle whose thermal efficiency is 40% receives 50 kJ by heat transfer from a hot reservoir at 600 K and rejects energy by heat transfer to a coldreservoir at temperature Tc. Determine the energy rejected in KJ
Answer #1
Effiiency=1-(Qc/Qh)=1-(Tc/Th)
0.4=1-(Qc/50)
Qc=30 kJ

1-Tc/600=0.4
Tc=360 K
Answer #2
nth=40 percent

1-(tc/th)=0.4

tc=360k


for reversible engine

1-qc/qh=0.4

qc=30kj
Answer #3

nth=40 percent

1-(tc/th)=0.4

tc=360k

by first law of thermodynamics

for reversible engine

1-qc/qh=0.4

qc=30kj